Twice Mike's age is the same as 3 times his brother's age. Togetehr they have been alive for 20 years. How old is each?

Answer by oberobic(2304)

2m = 3b
.
m + b = 20
so
m = 20-b
.
2(20-b) = 3b
.
40-2b = 3b
.
40 = 5b
.
5b = 40
.
b = 8
.
m = 20-8 = 12
.
2m = 2*12 = 24
3b = 3*8 = 24
OK
.
Answer: Mike is 12, and his brother is 8.
.
Done.